If the measure of each interior angle of a regular polygon is five times the measure of each exterior angle, how many sides does the polygon have ?
Answer by stanbon(75887)

If the measure of each interior angle of a regular polygon is five times the measure of each exterior angle, how many sides does the polygon have ?
---
ext + int = 180 degrees
int = 5*ext
--
ext + 5ext = 180
est = 30 degrees
----
sum of the exterior angles is 360
# of exterior = 360/30 = 12
---
# of sides is 12
===================
Cheers,
Stan H.
=============